opel do In randomized, double-blind clinical trials of a new vaccinerats were randomly divided into two groups. Subjects in g D. find the p score
E. Find the lower,upper bounds
Answer #1

Given,

\small p_1 = \frac{114}{665}=0.1714

\small p_2 = \frac{71}{546}=0.1300

we know that the test statistic is given by,

\small z = \frac{(p_1 - p_2)}{ \sqrt{\frac{p_1(1-p_1)}{n_1} +\frac{p_2(1-p_2)}{n_2} }}

\small z = \frac{(0.1714 - 0.1300)}{ \sqrt{\frac{0.1714(1-0.1714)}{665} +\frac{0.1300(1-0.1300)}{546} }}

\small z = 1.9922

The test statistic for the given hypothesis is,

1.99